OpenStudy (anonymous):

lim_{x rightarrow infty}(1-2x)e ^{-2x}

OpenStudy (anonymous):

Steps:\[\lim_{x \rightarrow \infty}(1-2x)(e^{-2x})\]\[\lim_{x \rightarrow \infty}\frac{1-2x}{e^{2x}}\] Applying L'hopital's rule: \[\lim_{x \rightarrow \infty} \frac{-2}{2.e^{2x}} = \lim_{x \rightarrow \infty}\frac{-2}{2.\infty} = \lim_{x \rightarrow \infty} \frac{-2}{\infty} = 0\] ^_^